OpenStudy (anonymous):

The general term of a sequence is given. Determine whether the given sequence is arithmetic, geometric, or neither. If the sequence is arithmetic, find the common difference; if it is geometric, find the common ratio. an = 4n - 2 A. arithmetic, d = -2 B. geometric, r = 4 C. arithmetic, d = 4 D. neither

OpenStudy (anonymous):

@cobra-strikes

OpenStudy (anonymous):

@dan815 @Hayleymeyer @cobra-strikes

OpenStudy (hayleymeyer):

oh wow o_o

OpenStudy (anonymous):

yep this is pre calc 12 grade math haha

OpenStudy (anonymous):

@ganeshie8

OpenStudy (hayleymeyer):

man, i get help you with that XD, are you good in spanish?

OpenStudy (anonymous):

yes

OpenStudy (hayleymeyer):

omg i love you! can you please help?!

OpenStudy (anonymous):

after my quiz breh

OpenStudy (anonymous):

@paki @Ashleyisakitty

OpenStudy (anonymous):

what is a1? what is a2? what is a3? what do the terms change by each time?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

when n=1 what is a1? plug in 1 and find out

OpenStudy (anonymous):

is it 2

OpenStudy (anonymous):

and how did you get that?

OpenStudy (anonymous):

i plug in one

OpenStudy (anonymous):

4(1) - 2 = 4 - 2 = 2. so what is a2?

OpenStudy (anonymous):

idk this is so confusing

OpenStudy (anonymous):

4(2) - 2 = ?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

-6? 4(2) -2 = 8 - 2 = ?

OpenStudy (welshfella):

the nth term of arithmetic sequence an = a1 + d(n-1) where d = common difference and a1 = first term

OpenStudy (anonymous):

now that makes more sence

OpenStudy (anonymous):

@cobra-strikes

OpenStudy (welshfella):

compare that general form to an = 4n - 2

OpenStudy (anonymous):

ok and what do you mean by that

OpenStudy (welshfella):

an = a1 + d(n -1) an =(a1 - d) + dn an = -2 + 4n

OpenStudy (welshfella):

comparing coefficients d = ?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

i guess that looks a bit confusing with the repeated d but its the last d which gives the common difference so d = 4

OpenStudy (welshfella):

a1 - d = -2 so if d = 4 a1 = 2 = first term

OpenStudy (welshfella):

the sequence is arithmetic 2 , 6 , 10 .....